Solve.
the length of a rectangle field is 6m less than twice the width. the perimeter is 312m. find the length of the feild.

Answer by checkley77(12844)

L=2W-6
2L+2W=312
2(2W-6)+2W=312
4W-12+2W=312
6W=312+12
6W=324
W=324/6
W=54 m. ans. for the width.
L=2*54-6=108-6=102 m. ans. for the length.
Proof:
2*102+2*54=312
204+108=312
312=312